### v1

# Hybrid Multi-stage Decoding for Few-shot NER with Entity-aware Contrastive Learning

Abstract
Few-shot named entity recognition can identify new types of named entities based on a few labeled examples. Previous methods employing token-level or span-level metric learning suffer from the computational burden and a large number of negative sample spans. In this paper, we propose the Hybrid Multi-stage Decoding for Few-shot NER with Entity-aware Contrastive Learning (MsFNER), which splits the general NER into two stages: entity-span detection and entity classification. There are 3 processes for introducing MsFNER: training, finetuning, and inference. In the training process, we train and get the best entity-span detection model and the entity classification model separately on the source domain using meta-learning, where we create a contrastive learning module to enhance entity representations for entity classification. During finetuning, we finetune the both models on the support dataset of target domain. In the inference process, for the unlabeled data, we first detect the entity-spans, then the entity-spans are jointly determined by the entity classification model and the KNN. We conduct experiments on the open FewNERD dataset and the results demonstrate the advance of MsFNER.
Hybrid Multi-stage Decoding for Few-shot NER with Entity-aware Contrastive Learning
Peipei Liu1,2, Gaosheng Wang1,2, Ying Tong3, Jian Liang3, Zhenquan Ding1thanks: The corresponding author, Hongsong Zhu1,2 1Institute of Information Engineering, Chinese Academy of Sciences 2School of Cyber Security, University of Chinese Academy of Sciences 3Jiangsu Provincial Department of Public Security, China {liupeipei, wanggaosheng, dingzhenquan, zhuhongsong}@iie.ac.cn
1 Introduction
Named Entity Recognition (NER) is a fundamental task in Natural Language Processing (NLP), involving the identification and categorization of text spans into predefined classes such as people, organizations, and locations Yadav and Bethard (2018); Li et al. (2022). Although traditional deep neural network architectures have achieved success in the fully supervised named entity recognition (NER) task with sufficient training data Lample et al. (2016); Ma and Hovy (2016), they are difficult to adapt to the dynamic nature of real-world applications, which require the model to quickly adjust itself to new data or environmental changes. In this case, researchers have proposed the few-shot named entity recognition (Few-shot NER) to explore entity recognition with limited labeled data Fritzler et al. (2019). Few-shot NER enables existing models to quickly transfer learned knowledge and adapt to new domains or entity classes.
Specifically, the Few-shot NER model is first trained on the source domain dataset $D_{source}=(S_{source},Q_{source})$ and then the trained model is transferred to new target domain dataset $D_{target}=(S_{target},Q_{target})$ to infer for $Q_{target}$ Snell et al. (2017). $S_{target}$ represents the support dataset comprising $N$ entity types ($N$-way), and each type is exemplified by $K$ annotated examples ($K$-shot). The $Q_{target}$ is the query dataset with the same entity types as the support dataset. Few-shot paradigm can offer a flexible and cost-effective solution to the adaptability challenge, making it a focal point of research to enhance the performance of NER systems in scenarios with limited labeled data or emerging entity types.
Currently, there are two mainstream researches for Few-shot NER: token-level Fritzler et al. (2019); Hou et al. (2020); Yang and Katiyar (2020); Das et al. (2022) and span-level metric learning methods Wang et al. (2022a); Yu et al. (2021); Ma et al. (2022). In token-level methods, each token is assigned an entity label based on its distances from the prototypes of entity classes or the support tokens. However, these approaches often have high computational costs and fail to maintain the semantic integrity of entity tokens, leading to increased susceptibility to interference from non-entity markers. On the other hand, although span-level methods can mitigate the partial issues associated with token-level approaches by evaluating entities as spans, all possible spans are enumerated would result in an N-square complexity and an increase in noise from a large number of negative samples.
Considering the challenges, we hope to solve the following problems: 1) To improve the Few-shot NER identification efficiency, how can we encourage the semantic divergence between entities and non-entities to determine effective entity spans? 2) To improve the entity span classification, how can we control and coordinate the semantic distance of different entity types to make the semantic representations of entities within the same types be proximate while those in disparate types be distant?
In this paper, we propose a hybrid multi-stage decoding approach for few-shot NER with contrastive learning (i.e., MsFNER). Specifically, MsFNER splits the general NER into two stages and this paper describes them by three processes: training process, finetuning process, and inference process. The training process consists of two objectives: training the best entity span detection model and the best entity classification model separately based on the supervised support dataset of source domain. In this process, we regard the training task as a meta-learning task and train entity detection model by employing Model-Agnostic Meta-Learning(MAML)Finn et al. (2017), and train entity classification model through the entity-aware contrastive learning and MAML-ProtoNetMa et al. (2022). In the finetuning process, we finetune the trained entity detection model and entity classification model on the support dataset of target domain with the same pattern as training process. In the inference process, we deploy the KNN and two finetuned models on the query dataset of target domain to predict entity spans and their types, and there are four phases for prediction. In the first phase, we build a key-value datastore $\mathcal{D}_{knn}$ on the support dataset $S_{target}$, where key is the entity representation and value is the corresponding label. Subsequently, we apply the entity detection model to get entity spans. With the entity detection model, we could reduce the impact of negative samples and reduce computational complexity. After that, we pass the representations of detected entity spans into the entity classification model to produce their predicted type distribution $p_{soft}$ while we also feed the representations into KNN to obtain the predicted results $p_{knn}$ based on the $\mathcal{D}_{knn}$. In the final phase, we combine $p_{soft}$ and $p_{knn}$ to obtain final types for prediction entities.
We summarize our contributions as follows:
  • We propose a hybrid multi-stage decoding for few-shot NER with entity-aware contrastive learning, in which we first detect entity spans to improve efficiency and then employ contrastive learning or KNN to augment entity classification performance.
  • Experimental results show that our model achieves new SOTA performance compared with previous works. We also evaluate the few-shot NER task on LLM ChatGPT and the experimental results show that our model outperforms ChatGPT in terms of performance and efficiency.
2 Methodology
In this section, we will introduce the details of MsFNER. The following content is arranged according to the three stages of the model.
2.1 Training Process
In this process, we introduce the separate process of training and finding the optimal entity span detection model and entity classification model.
Figure 1: The schematic diagram of the training process.
2.1.1 Entity Span Detection (ESD)
[…]
For a given sentence $x$ with $n$ tokens $x$=$(x_{1},x_{2},...,x_{n})$ in $D_{source}$, we first encode it using the pre-trained language model BERT. After that, we can obtain the contextualized representations $h=(h_{1},h_{2},...,h_{n})$ for all tokens.
The sequence token representations are then sent into a CRF to detect entity spans. Finally, the training loss in $D_{source}$ can be depicted as:
$L_{ESD}=-\sum logP(y|x)$ (1)
$P(y|x)=\frac{\prod_{i=1}^{|x|}\varphi_{i}(y_{i-1},y_{i},x)}{\sum_{y^{{}^{\prime}}}\prod_{i=1}^{|x|}\varphi_{i}(y_{i-1}^{{}^{\prime}},y_{i}^{{}^{\prime}},x)}$ (2)
where $\varphi_{i}(y_{i-1},y_{i},x)$ and $\varphi_{i}(y_{i-1}^{{}^{\prime}},y_{i}^{{}^{\prime}},x)$ are potential functions.
We regard the training process as a meta-learning task and divide the $D_{source}$ into many individual meta learning tasks. Then, we train the best entity-span detection model by MAML.
2.1.2 Entity Classification
For an entity $e_{k}=(x_{f},...,x_{f+l})$, where $f$ is the first token index and $f+l$ is the last token index in the sequence, we employ the max-pooling to get its representation when $l>=1$ :
$\hat{e}_{k}=max(h_{f},...,h_{f+l})$ (3)
Contrastive learning could enhance the consistency between entities within the same types and widen the distance between entities belonging to different types. As $e_{k}$ is the sample with certain type from the supervised $D_{source}$, we consider using the supervised contrastive learning to deepen the distinctiveness of entity types and augment representations of entities for entity classification improvement.
[…]
$L_{CL}=\sum_{j}-\frac{1}{|P(j)|}\sum_{p\in P(j)}log\frac{exp(sim(z^{j},z^{p})/\tau)}{\sum_{a\in A(j)}exp(sim(z^{j},z^{a})/\tau)}$ (4)
where P(j) is the positive set whose entities are from the same type with $e_{j}$, $z^{j}$ is the result of transforming $\hat{e}_{j}$ through a projection network MLP, A(j) is the set containing all N entities except for $e_{j}$, sim(·,·) denotes the KL-divergence for $K$-shot ($K$>1) but the squared euclidean distance for $K$-shot ($K$=1) referring to Das et al. (2022), and $\tau$ is a temperature hyperparameter.
After pulling entities of the same type together and entities of different types farther apart in the semantic space with contrastive learning, we discard the projection network at classification time following the previous works of contrastive learning Chen et al. (2020); Khosla et al. (2020); Liu et al. (2023). As a consequence, we use the entity representation $\hat{e}_{k}$ to compute the prototype $c_{t}(\mathcal{S})$ of $t_{th}$ type Snell et al. (2017), and predict the type distribution of $e_{k}$ by:
$p_{soft}(e_{k})=\frac{\exp\left(-d\left(c_{t}(\mathcal{S}),\hat{e}_{k}\right)\right)}{\sum_{ii\in\phi}\exp\left(-d\left(c_{ii}(\mathcal{S}),\hat{e}_{k}\right)\right)}$ (5)
[…]
2.2 Finetuning Process
Model finetuning aims to enable the model to adapt to new knowledge types during the domain transfer process. We finetune the trained models for entity detection and entity classification on the support dataset of the target domain $S_{target}$, using the same method as training process.
2.3 Inference Process
Figure 2: The schematic diagram of the inference process for entity classification.
In the inference process, we test the finetuned model on query dataset $Q_{target}$, and we will present this stage by four parts:
(1) We compute the representation $\hat{e}_{k}$ of each true entity in the support dataset $S_{target}$ by Eq.3 and build a key-value datastore $\mathcal{D}_{knn}$ where key is the entity representation and value is entity type.
(2) We input the query sentence $x^{\prime}$ into the finetuned best model for entity span detection (ESD) to get all entities. Specifically, we apply the Viterbi algorithm Forney (1973) for decoding, and derive the entities from the maximum sequence probability of Eq.2:
$y^{*}=argmax_{y\in\mathbb{Y}}P(y|x^{\prime})$ (7)
where $\mathbb{Y}$ is a set of all possible label sequences.
Models FewNERD-INTRA FewNERD-INTER
1-2 shot 5-10 shot 1-2 shot 5-10 shot
5 way 10 way 5 way 10 way 5 way 10 way 5 way 10 way
ProtoBERT† 23.45±0.92 19.76±0.59 41.93±0.55 34.61±0.59 44.44±0.11 39.09±0.87 58.80±1.42 53.97±0.38
NNShot† 31.01±1.21 21.88±0.23 35.74±2.36 27.67±1.06 54.29±0.40 46.98±1.96 50.56±3.33 50.00±0.36
StructShot† 35.92±0.69 25.38±0.84 38.83±1.72 26.39±2.59 57.33±0.53 49.46±0.53 57.16±2.09 49.39±1.77
CONTAINER† 40.436 33.84 53.70 47.49 55.95 48.35 61.83 57.12
ESD† 41.44±1.16 32.29±1.10 50.68±0.94 42.92±0.75 66.46±0.49 59.95±0.69 74.14±0.80 67.91±1.41
MAML-ProtoNet† 52.04±0.44 43.50±0.59 63.23±0.45 56.84±0.14 68.77±0.24 63.26±0.40 71.62±0.16 68.32±0.10
ChatGPT‡ 61.86 55.61 64.17 54.79 61.05 57.94 64.34 58.58
MsFNER (Ours) 54.25±0.34 46.69±0.5 66.57±0.29 58.70±1.39 72.91±0.34 66.34±0.65 78.41±0.30 72.06±0.11
w/o cl* 65.72±0.33 57.33±0.08 77.84±0.42 71.23±0.47
w/o KNN 53.86±0.47 46.24±0.29 66.04±0.65 57.56±0.41 72.75±0.29 65.38±0.72 77.99±0.25 71.92±0.14
Table 1: F1 scores with standard deviations on FewNERD. †denotes the results reported in Ma et al. (2022). ‡are the results we produce by LLM. * means the abbreviation of ‘contrastive learning’. The best results are in bold.
(3) Firstly, we compute the representation $\hat{e}^{\prime}_{k}$ of each detected entity ${e}^{\prime}_{k}$ by Eq.3. Then, we input $\hat{e}^{\prime}_{k}$ into two independent modules: on one hand, we feed the $\hat{e}^{\prime}_{k}$ into KNN to obtain the predicted result $p_{knn}$ based on the $\mathcal{D}_{knn}$ Wang et al. (2022b), and on the other hand we pass the $\hat{e}^{\prime}_{k}$ into the finetuned best model for entity classification to produce its predicted type distribution $p_{soft}$ by Eq.5.
(4) The final prediction type of the detected entity ${e}^{\prime}_{k}$ is jointly determined by KNN and the entity classification model:
$p(y|e^{\prime}_{k})=\lambda p_{knn}(y|e^{\prime}_{k})+(1-\lambda)p_{soft}(y|e^{\prime}_{k})$ (8)
where the $\lambda$ is a hyper-parameter that makes a balance between KNN and softmax.
3 Experiments
3.1 Experiments Setup
The details can be found at A.1.
3.2 Main results
Table 1 illustrates the main results of different methods on FewNERD dataset. The results demonstrate that MsFNER significantly outperforms all the previous state-of-the-art approaches. MsFNER attains average F1 improvements of 2.65 and 4.44 on INTRA and INTER, respectively, compared to the previous best method, MAML-ProtoNet. In the 5-way 5-10 shot setting on the INTER dataset, we achieve the most substantial improvement over the MAML-ProtoNet method, reaching 7.79 points. Additionally, we perform the few-shot task employing ChatGPT, where MsFNER surpasses ChatGPT in F1 performance by an average score of 8.465 in the 5-10 shot setting. Furthermore, MsFNER surpasses ChatGPT by an average of 10.13 percentage points in the 1-2 shot setting in the INTER dataset. Overall, regardless of whether it is ChatGPT or small models, increasing the number of $K$-shot yields superior performance, while an increase in the number of $N$-way results in worse performance.
3.3 Ablation Study
To investigate the contributions of different modules of MsFNER, we conduct the ablation study by removing each of them at a time to observe the performance of our model. The results are shown at the bottom of Table 1. Firstly, we remove the contrastive learning module. In the 1 shot experiment of this setting, KNN is unnecessary as there is only one sample in the support dataset. We can see that the removal of contrastive learning results in a decrease of 0.905 average score, which indicates the significant impact and necessity of contrastive learning for enhancing representation. Secondly, we remove the KNN module and reserve the contrastive learning module. The removal of KNN drops 0.523 average score. We can see that the impact of contrastive learning and KNN increases with the number of $N$-way and $K$-shot. Removing either of them will lead to a significant decrease in the performance of MsFNER.
3.4 Model Efficiency
To evaluate the efficiency of MsFNER, we compute the average inference time of MsFNER and ChatGPT in different settings. More details can be found at A.3.
4 Conclusion
In this paper, we propose the MsFNER, aiming to improve the performance and efficiency of few-shot NER. The MsFNER can be achieved by three stages: training, finetuning and inferring. We evaluate MsFNER on the open FewNERD dataset, and results show the advance of MsFNER compared with previous few-shot NER methods and LLM.
5 Limitations
  • In practical, the effectiveness of Few-shot NER is seriously affected due to the data sparsity and data imbalance.
  • The abundant external knowledge is very helpful for few-shot NER, which is worth exploring and studying.
6 Ethical Considerations
All the data we get is the open source, and there is no theft. We guarantee the originality, research ethics, societal impact and reproducibility of our work.
Appendix A Appendix
A.1 Experiments Setup
We conduct experiments on the Tesla V100 GPU.
A.1.1 Datasets
We conduct experiments on the FewNERD dataset Ding et al. (2021) which is annotated with 8 coarse-grained and 66 fine-grained entity types. We evaluate our method on the two settings of FewNERD: 1) FewNERD-INTRA, which pertains to a scenario wherein entities within the training set (source domain), validation set, and test set (target domain) are exclusively associated with distinct coarse-grained types. 2) FewNERD-INTER, where only fine-grained entity types that do not intersect in different sets. Furthermore, we follow the settings of ESD Wang et al. (2022a) and adopt $N$-way $K\sim 2K$-shot sampling method to construct tasks.
A.1.2 Parameter Settings
We use grid search for hyperparameter settings and we train our model for 1,000 steps and choose the best model on validation dataset for testing. We use the pretrained language model uncased BERT-base as our encoder. In the entity span detection module, we adopt the BIOES tags. We use the AdamW optimizer with a learning rate of 3e-5 and 0.01 linear warmup steps. The $K$ in KNN is set to 10 and the max-loss coefficient $\lambda$ of KNN is 0.1 when adding KNN. The batch size is set to 32, the max sequence length is set to 128 and we use a dropout probability of 0.2.
Models FewNERD-INTRA FewNERD-INTER
1-2 shot 5-10 shot 1-2 shot 5-10 shot
5 way 10 way 5 way 10 way 5 way 10 way 5 way 10 way
MsFNER 0.3442 0.3808 0.5616 1.0546 0.3482 0.3806 0.6708 1.2266
ChatGPT 4.5581 5.4897 8.1842 9.3481 3.2633 8.0579 4.7755 5.3409
Table 2: The comparison of inference time between MsFNER and ChatGPT.
A.1.3 Baselines
We compare MsFNER with popular baselines as follows:
  • ProtoBERT Fritzler et al. (2019) employs the prediction of query labels by leveraging the similarity between the BERT hidden states derived from the support set and the query tokens.
  • CONTAINER Das et al. (2022) employs token-level contrastive learning to train BERT as the token embedding function. Subsequently, BERT undergoes finetuning on the support set, followed by the application of a nearest neighbor method during the inference phase.
  • NNShot Yang and Katiyar (2020) bears resemblance to ProtoBERT in its methodology, as it conducts predictions utilizing nearest neighbor algorithms.
  • StructShot Yang and Katiyar (2020) incorporates an augmented Viterbi decoder into the inference process atop NNShot’s framework.
  • ESD Wang et al. (2022a) formulates few-shot sequence labeling as a span-level matching problem and decomposes it into span-related procedures.
  • MAML-ProtoNet Ma et al. (2022) proposes a decomposed meta-learning approach for few-shot NER, which tackles few-shot span detection and entity typing using MAML to facilitate rapid adaptation to novel entity classes and proposes MAML-ProtoNet for improved representation of entity spans.
  • ChatGPT chatgpt (2023) is an advanced neural language model developed by OpenAI, which is trained on diverse corpora of text and could generate human-like text.
A.2 ChatGPT Prompt Template
In this section, we introduce the ChatGPT few-shot NER prompt template that we use. The prompt is composed of 3 parts: task description, few-shot cases, and input queries.
In the task description, we initially elucidate the few-shot Named Entity Recognition (NER) task required from ChatGPT. Subsequently, we provide ChatGPT with explicit instructions for implementing this task, including delineation of the entity types involved. To facilitate easier interpretation of model output, we utilize a structured output JSON format, a domain in which ChatGPT exhibits proficiency. The template is as follows:
Your task is to perform fewshot Named Entity Recognition. You could perform the task by the following actions:
  1. 1.
    Do named entity recognition task to recognize the entities. Entity type is defined as \$ent_list
  2. 2.
    Check if the entity type is in the definition list.
  3. 3.
    Output a JSON object that contains the following keys: text, entity_list. Note that each item in entity_list is a dictionary with "entity" and "type" as keys.
Here are the given few-shot cases: ‘‘‘\$case_input‘‘‘
Output: \$case_output
Recognize the entities in the text delimited by triple backticks.: ‘‘‘\$input_text‘‘‘
In the template above, the \$ent_list denotes the list of entity types corresponding to $N$-way. The \$case_input comprises the sentences in the support set corresponding to $K$-shot, while the \$case_output represents the labels in the support set. Since ChatGPT is better at structured output, we use JSON format for output, with each output format being: {"entity": entity text ; "type": entity type}. The output will be merged into a list. The \$input_text denotes the query data. When we input this prompt to ChatGPT, it will return a list of relevant entity-type dictionaries.
A.3 Model Efficiency
To evaluate the efficiency of MsFNER, we compute the average inference time of MsFNER and ChatGPT in different settings. And the results are shown in Table 2, where we can see that MsFNER is on the millisecond level while the ChatGPT is on the second level. The time consumption is related to the amount of input data. In addition, the time consumption of ChatGPT will increase due to network impact.
A.4 training process
As the Figure 3 shows, the training loss on training dataset is smoothly declining within the same source types. But when finetuning on the validation set, there are some fluctuations due to the cross domain transfer of knowledge. And with the training time increasing, the overfitting is occurred slightly on validation dataset because of few supervised samples.
Figure 3: The performance of training process on FewNERD 5-way 5~10-shot. (a) is the evaluation F1 score when finetuning on the validation dataset. (b) is the training loss on train dataset.

### v2

@@
 During finetuning, we finetune the both models on the support dataset of target domain.
-In the inference process, for the unlabeled data, we first detect the entity-spans, then the entity-spans are jointly determined by the entity classification model and the KNN.
-We conduct experiments on the open FewNERD dataset and the results demonstrate the advance of MsFNER.
-Hybrid Multi-stage Decoding for Few-shot NER with Entity-aware Contrastive Learning
-Peipei Liu1,2, Gaosheng Wang1,2, Ying Tong3, Jian Liang3, Zhenquan Ding1thanks: The corresponding author, Hongsong Zhu1,2 1Institute of Information Engineering, Chinese Academy of Sciences 2School of Cyber Security, University of Chinese Academy of Sciences 3Jiangsu Provincial Department of Public Security, China {liupeipei, wanggaosheng, dingzhenquan, zhuhongsong}@iie.ac.cn
-1 Introduction
-Named Entity Recognition (NER) is a fundamental task in Natural Language Processing (NLP), involving the identification and categorization of text spans into predefined classes such as people, organizations, and locations Yadav and Bethard (2018); Li et al.
-(2022).
-Although traditional deep neural network architectures have achieved success in the fully supervised named entity recognition (NER) task with sufficient training data Lample et al.
-(2016); Ma and Hovy (2016), they are difficult to adapt to the dynamic nature of real-world applications, which require the model to quickly adjust itself to new data or environmental changes.
-In this case, researchers have proposed the few-shot named entity recognition (Few-shot NER) to explore entity recognition with limited labeled data Fritzler et al.
-(2019).
+In the inference process, for the unlabeled query data, we first detect the entity-spans, then the entity-spans are jointly determined by the entity classification model and the KNN.
+We conduct experiments on the open FewNERD dataset and FewAPTER dataset, the results demonstrate the advance of MsFNER.
+I Introduction
+Named Entity Recognition (NER) is a fundamental task in Natural Language Processing (NLP), involving the identification and categorization of text spans into predefined classes such as people, organizations, and locations [1, 2].
+Although traditional deep neural network architectures have achieved success in the fully supervised named entity recognition task with sufficient training data [3, 4], they are difficult to adapt to the dynamic nature of real-world applications, which require the model to quickly adjust itself to new data or environmental changes.
+In this case, researchers have proposed the few-shot named entity recognition (few-shot NER) to explore entity recognition with limited labeled data [5].
 Few-shot NER enables existing models to quickly transfer learned knowledge and adapt to new domains or entity classes.
-Specifically, the Few-shot NER model is first trained on the source domain dataset $D_{source}=(S_{source},Q_{source})$ and then the trained model is transferred to new target domain dataset $D_{target}=(S_{target},Q_{target})$ to infer for $Q_{target}$ Snell et al.
-(2017).
+Specifically, the few-shot NER model is first trained on the source domain dataset $D_{source}=(S_{source},Q_{source})$ and then the trained model is transferred to new target domain dataset $D_{target}=(S_{target},Q_{target})$ to infer for $Q_{target}$ [6].
 $S_{target}$ represents the support dataset comprising $N$ entity types ($N$-way), and each type is exemplified by $K$ annotated examples ($K$-shot).
@@
 Few-shot paradigm can offer a flexible and cost-effective solution to the adaptability challenge, making it a focal point of research to enhance the performance of NER systems in scenarios with limited labeled data or emerging entity types.
-Currently, there are two mainstream researches for Few-shot NER: token-level Fritzler et al.
-(2019); Hou et al.
-(2020); Yang and Katiyar (2020); Das et al.
-(2022) and span-level metric learning methods Wang et al.
-(2022a); Yu et al.
-(2021); Ma et al.
-(2022).
+Currently, there are two mainstream researches for few-shot NER: token-level [5, 7, 8, 9] and span-level metric learning methods [10, 11, 12].
 In token-level methods, each token is assigned an entity label based on its distances from the prototypes of entity classes or the support tokens.
@@
 The training process consists of two objectives: training the best entity span detection model and the best entity classification model separately based on the supervised support dataset of source domain.
-In this process, we regard the training task as a meta-learning task and train entity detection model by employing Model-Agnostic Meta-Learning(MAML)Finn et al.
-(2017), and train entity classification model through the entity-aware contrastive learning and MAML-ProtoNetMa et al.
-(2022).
+In this process, we regard the training task as a meta-learning task and train entity detection model by employing Model-Agnostic Meta-Learning(MAML)[13], and train entity classification model through the entity-aware contrastive learning and MAML-ProtoNet[12].
 In the finetuning process, we finetune the trained entity detection model and entity classification model on the support dataset of target domain with the same pattern as training process.
@@
 After that, we pass the representations of detected entity spans into the entity classification model to produce their predicted type distribution $p_{soft}$ while we also feed the representations into KNN to obtain the predicted results $p_{knn}$ based on the $\mathcal{D}_{knn}$.
-In the final phase, we combine $p_{soft}$ and $p_{knn}$ to obtain final types for prediction entities.
+In the final phase, we combine $p_{soft}$ and $p_{knn}$ for predicting entity types.
 We summarize our contributions as follows:
@@
   • Experimental results show that our model achieves new SOTA performance compared with previous works.
-We also evaluate the few-shot NER task on LLM ChatGPT and the experimental results show that our model outperforms ChatGPT in terms of performance and efficiency.
-2 Methodology
+We also evaluate the few-shot NER task on LLM ChatGPT, and MsFNER can have the comparable comprehensive results with ChatGPT in terms of performance and efficiency.
+II Methodology
 In this section, we will introduce the details of MsFNER. The following content is arranged according to the three stages of the model.
-2.1 Training Process
+Figure 1: The schematic diagram of the training process.
+II-A Training Process
 In this process, we introduce the separate process of training and finding the optimal entity span detection model and entity classification model.
-Figure 1: The schematic diagram of the training process.
-2.1.1 Entity Span Detection (ESD)
+Entity Span Detection (ESD)
+II-A1 ESD Method
 […]
-For a given sentence $x$ with $n$ tokens $x$=$(x_{1},x_{2},...,x_{n})$ in $D_{source}$, we first encode it using the pre-trained language model BERT.
-After that, we can obtain the contextualized representations $h=(h_{1},h_{2},...,h_{n})$ for all tokens.
+[…]
 The sequence token representations are then sent into a CRF to detect entity spans. Finally, the training loss in $D_{source}$ can be depicted as:
@@
 $P(y|x)=\frac{\prod_{i=1}^{|x|}\varphi_{i}(y_{i-1},y_{i},x)}{\sum_{y^{{}^{\prime}}}\prod_{i=1}^{|x|}\varphi_{i}(y_{i-1}^{{}^{\prime}},y_{i}^{{}^{\prime}},x)}$ (2)
-where $\varphi_{i}(y_{i-1},y_{i},x)$ and $\varphi_{i}(y_{i-1}^{{}^{\prime}},y_{i}^{{}^{\prime}},x)$ are potential functions.
+where $\varphi_{i}(y_{i-1},y_{i},x)$ and $\varphi_{i}(y_{i-1}^{{}^{\prime}},y_{i}^{{}^{\prime}},x)$ are potential functions, $y^{\prime}$ is the possible label sequence.
+II-A2 Training
 We regard the training process as a meta-learning task and divide the $D_{source}$ into many individual meta learning tasks. Then, we train the best entity-span detection model by MAML .
-2.1.2 Entity Classification
+Specifically, the model first randomly samples a batch of task data $\mathcal{T}=\{(S_{source}^{(i)},Q_{source}^{(i)})\}_{i=1}^{bz}$ from the source domain data $D_{source}$, where the batch size is $bz$.
+Next, the model is trained separately on $\mathcal{T}_{s}=\{S_{source}^{(i)}\}_{i=1}^{bz}$ and $\mathcal{T}_{q}=\{Q_{source}^{(i)}\}_{i=1}^{bz}$, which we refer to as inner-loop update (Task-level adaptation) and outer-loop update (Meta-optimization), respectively.
+An overview of the training process is presented in Algorithm 1.
+In the inner-loop update, for each learning episode in meta-learning, the model performs one gradient update on the task data $S_{source}^{(i)}$ to update the parameters:
+$\theta_{i}^{\prime}=\theta-\alpha\nabla_{\theta}{L}_{\mathit{ESD}}\left(S_{source}^{(i)};{\theta}\right)$ (3)
+where $\theta=\{\theta^{encoder},\theta^{crf}\}$ represents the model’s learnable parameters, $\theta_{i}^{\prime}$ is the updated parameter after one gradient update on the task data $S_{source}^{(i)}$, and $\alpha$ is the learning rate used to minimize the loss function (Equation (1)).
+After the inner-loop update, the model will fully utilize the labeled query set $\mathcal{T}_{q}$ from the source domain for the outer-loop update.
+This update differs from the inner-loop update’s model optimization approach (i.e., using the support set of one task in each learning episode for parameter optimization).
+Instead, it follows the conventional neural network training method, where the loss from all the training data in the batch is optimized together.
+The main purpose of using both inner and outer-loop update is to enhance the model’s generalization ability on new task data and prevent overfitting to the support set data.
+The optimization objective of the outer-loop update is:
+$\min_{\theta}\sum\nolimits_{\mathcal{T}_{q}}{L}_{\mathit{ESD}}\left(Q_{source}^{(i)};\theta_{i}^{\prime}\right)$ (4)
+Although the loss calculation at this point involves $\theta_{i}^{\prime}$, the model’s goal remains to update and optimize $\theta$.
+The outer-loop update optimization is performed through Stochastic Gradient Descent (SGD), and therefore, the model parameter update process can be expressed as:
+$\theta=\theta-\beta\sum\nolimits_{\mathcal{T}_{q}}\nabla_{\theta_{i}^{\prime}}{L}_{\mathit{ESD}}\left(Q_{source}^{(i)};\theta_{i}^{\prime}\right)$ (5)
+where $\beta$ is the learning rate for SGD.
+Algorithm 1 Training Process of Entity Span Detection Model Based on MAML
+[…]
+[…]
+3:Randomly initialize $\theta$
+4:while not done do
+5:  Randomly select a batch of task data $\mathcal{T}$, where $\mathcal{T}\sim D_{source}$
+6:  for all $\mathcal{T}$ do
+7:   Select one task data $(S_{source}^{(i)},Q_{source}^{(i)})$ from $\mathcal{T}$
+8:   Compute the gradient $\nabla_{\theta}{L}_{\mathit{ESD}}\left(S_{source}^{(i)};{\theta}\right)$ based on the support set $S_{source}^{(i)}$
+9:   Perform a gradient update: $\theta_{i}^{\prime}=\theta-\alpha\nabla_{\theta}{L}_{\mathit{ESD}}\left(S_{source}^{(i)};{\theta}\right)$
+10:   Compute the gradient $\nabla_{\theta_{i}^{\prime}}{L}_{\mathit{ESD}}\left(Q_{source}^{(i)};\theta_{i}^{\prime}\right)$ based on the query set $Q_{source}^{(i)}$, and store the result
+11:  end for
+12:  Outer-loop update: $\theta=\theta-\beta\sum_{\mathcal{T}_{q}}\nabla_{\theta_{i}^{\prime}}{L}_{\mathit{ESD}}\left(Q_{source}^{(i)};\theta_{i}^{\prime}\right)$
+13:end while
+Entity Classification (EC)
+II-A1 EC Method
 For an entity $e_{k}=(x_{f},...,x_{f+l})$, where $f$ is the first token index and $f+l$ is the last token index in the sequence, we employ the max-pooling to get its representation when $l>=1$ :
@@
 As $e_{k}$ is the sample with certain type from the supervised $D_{source}$, we consider using the supervised contrastive learning to deepen the distinctiveness of entity types and augment representations of entities for entity classification improvement.
-[…]
+In details, given a batch with N entities and the anchor index j $\in$ {1,2,…,N}, an entity-aware contrastive loss can be defined as follows:
 $L_{CL}=\sum_{j}-\frac{1}{|P(j)|}\sum_{p\in P(j)}log\frac{exp(sim(z^{j},z^{p})/\tau)}{\sum_{a\in A(j)}exp(sim(z^{j},z^{a})/\tau)}$ (7)
-where P(j) is the positive set whose entities are from the same type with $e_{j}$, $z^{j}$ is the result of transforming $\hat{e}_{j}$ through a projection network MLP, A(j) is the set containing all N entities except for $e_{j}$, sim(·,·) denotes the KL-divergence for $K$-shot ($K$>1) but the squared euclidean distance for $K$-shot ($K$=1) referring to Das et al.
-(2022), and $\tau$ is a temperature hyperparameter.
-After pulling entities of the same type together and entities of different types farther apart in the semantic space with contrastive learning, we discard the projection network at classification time following the previous works of contrastive learning Chen et al.
-(2020); Khosla et al.
-(2020); Liu et al.
-(2023).
-As a consequence, we use the entity representation $\hat{e}_{k}$ to compute the prototype $c_{t}(\mathcal{S})$ of $t_{th}$ type Snell et al.
-(2017), and predict the type distribution of $e_{k}$ by:
-$p_{soft}(e_{k})=\frac{\exp\left(-d\left(c_{t}(\mathcal{S}),\hat{e}_{k}\right)\right)}{\sum_{ii\in\phi}\exp\left(-d\left(c_{ii}(\mathcal{S}),\hat{e}_{k}\right)\right)}$ (5)
+where P(j) is the positive set whose entities are from the same type with $e_{j}$, $z^{j}$ is the result of transforming $\hat{e}_{j}$ through a projection network MLP, A(j) is the set containing all N entities except for $e_{j}$, sim(·,·) denotes the KL-divergence for $K$-shot (K>K>1) but the squared euclidean distance for $K$-shot (KK=1) referring to [9], and $\tau$ is a temperature hyperparameter.
+After pulling entities of the same type together and entities of different types farther apart in the semantic space with contrastive learning, we discard the projection network at classification time following the previous works of contrastive learning [14, 15, 16].
+Subsequently, prototypes for different types of entities are constructed based on the enhanced entity features.
+Let $\mathcal{S}_{t}$ denote the set of entities belonging to the $t$-th type in the support dataset $S_{source}^{(i)}$ of a task during training.
+The prototype for the $t$-th type is obtained by calculating the average of all entity features in this type.
+$c_{t}(\mathcal{S})=\frac{1}{\left|S_{t}\right|}\sum_{{e}_{m}\in S_{t}}\hat{e}_{m}$ (8)
+As a consequence, we can predict the type distribution of $e_{k}$ by calculating the distance between its entity representation $\hat{e}_{k}$ and all type prototypes:
+$p_{soft}(e_{k})=\frac{\exp\left(-d\left(c_{t}(\mathcal{S}),\hat{e}_{k}\right)\right)}{\sum_{ii=1}^{|\phi|}\exp\left(-d\left(c_{ii}(\mathcal{S}),\hat{e}_{k}\right)\right)},t\in\{1,..,|\phi|\}$ (9)
+where $\phi$ is the entity type set and $|\phi|$=NN, $d(\cdot,\cdot)$ is the distance function. So, the entity classification loss is formalized as:
+$L_{EC}=-\sum_{k=1}^{NK}y_{k}log(p_{soft}({e}_{k}))$ (10)
+where $y_{k}$ is the truth type of ${e}_{k}$, $L_{\mathit{EC}}$ is the loss which we optimize.
+II-A2 Training
+The training of the model in this stage (detailed in Algorithm 2) adopts the MAML-ProtoNet approach proposed by Ma et al.
+[12], which combines the MAML method with prototype network techniques.
+Essentially, the key difference from MAML lies in treating the prototypes of entity types in the support set as type labels for entities of the same type during training.
+In the inner-loop update, the model first uses Equation (9) to predict the types for all true entity spans in the support set $S_{source}^{(i)}$ of each task.
+Then, the entity classification loss is computed based on $c^{N}_{k=1}(\mathcal{S})$ and Equation (10).
+Finally, gradient computation is performed using Equation (3) to optimize the parameters.
+In the outer-loop update, the training process is similar to entity span detection.
+The model first uses Equation (9) to predict the types for all true entity spans in the query set $Q_{source}^{(i)}$ of each task.
+Then, based on $c^{N}_{k=1}(\mathcal{S})$ and Equation (10), the entity classification loss is computed.
+Subsequently, as shown in Equation (4), the losses across all tasks are accumulated.
+Finally, gradient computation is performed using Equation (5) to optimize the parameters.
+Algorithm 2 Training Process of Entity Classification Model Based on MAML-ProtoNet Algorithm
 […]
 […]
-2.2 Finetuning Process
+3:Randomly initialize $\theta$
+4:while not done do
+5:  Randomly select a batch of data $\mathcal{T}$, $\mathcal{T}\sim D_{source}$
+6:  for all $\mathcal{T}$ do
+7:   Select a task $(S_{source}^{(i)},Q_{source}^{(i)})$ from $\mathcal{T}$
+8:   Obtain $c^{N}_{k=1}(\mathcal{S})$ by computing prototype vectors on $S_{source}^{(i)}$ using Equation (8)
+9:   Based on $c^{N}_{k=1}(\mathcal{S})$ and the true entity spans in the support set $S_{source}^{(i)}$, compute $\nabla_{\theta}{L}_{\mathit{EC}}\left(S_{source}^{(i)};{\theta}\right)$ using Equations (9, 10)
+10:   Perform gradient update: $\theta_{i}^{\prime}=\theta-\alpha\nabla_{\theta}{L}_{\mathit{EC}}\left(S_{source}^{(i)};{\theta}\right)$
+11:   Based on $c^{N}_{k=1}(\mathcal{S})$ and the true entity spans in the query set $Q_{source}^{(i)}$, compute $\nabla_{\theta_{i}^{\prime}}{L}_{\mathit{EC}}\left(Q_{source}^{(i)};\theta_{i}^{\prime}\right)$ using Equations (9, 10) and save the results
+12:  end for
+13:  Outer-loop update: $\theta=\theta-\beta\sum\nolimits_{\mathcal{T}_{q}}\nabla_{\theta_{i}^{\prime}}{L}_{\mathit{EC}}\left(Q_{source}^{(i)};\theta_{i}^{\prime}\right)$
+14:end while
+II-B Finetuning Process
 Model finetuning aims to enable the model to adapt to new knowledge types during the domain transfer process.
-We finetune the trained models for entity detection and entity classification on the support dataset of the target domain $S_{target}$, using the same method as training process.
-2.3 Inference Process
-Figure 2: The schematic diagram of the inference process for entity classification.
+We finetune the trained models for entity detection and entity classification on the support dataset of the target domain $S_{target}$, using the same pattern as training process.
+II-C Inference Process
 In the inference process, we test the finetuned model on query dataset $Q_{target}$, and we will present this stage by four parts:
@@
 (2) We input the query sentence $x^{\prime}$ into the finetuned best model for entity span detection (ESD) to get all entities.
-Specifically, we apply the Viterbi algorithm Forney (1973) for decoding, and derive the entities from the maximum sequence probability of Eq.2:
+Specifically, we apply the Viterbi algorithm [17] for decoding, and derive the entities from the maximum sequence probability of Eq.2:
 $y^{*}=argmax_{y\in\mathbb{Y}}P(y|x^{\prime})$ (11)
 where $\mathbb{Y}$ is a set of all possible label sequences.
+TABLE I: F1 scores with standard deviations(%) on FewNERD.
+†denotes the results reported in [12] and [18].
+‡are the results we produce by LLM.
+* means the abbreviation of ‘contrastive learning’.
+The best results are in bold.
 Models FewNERD-INTRA FewNERD-INTER
@@
 MAML-ProtoNet† 52.04±0.44 43.50±0.59 63.23±0.45 56.84±0.14 68.77±0.24 63.26±0.40 71.62±0.16 68.32±0.10
+BDCP† 52.63 - - 57.46 69.59 - - 68.87
 ChatGPT‡ 61.86 55.61 64.17 54.79 61.05 57.94 64.34 58.58
@@
 w/o KNN 53.86±0.47 46.24±0.29 66.04±0.65 57.56±0.41 72.75±0.29 65.38±0.72 77.99±0.25 71.92±0.14
-Table 1: F1 scores with standard deviations on FewNERD.
-†denotes the results reported in Ma et al.
-(2022).
-‡are the results we produce by LLM.
-* means the abbreviation of ‘contrastive learning’.
-The best results are in bold.
+TABLE II: The comparison of inference time between MsFNER and ChatGPT.
+Models FewNERD-INTRA FewNERD-INTER FewAPTER
+1-2 shot 5-10 shot 1-2 shot 5-10 shot 1 shot 3 shot
+5 way 10 way 5 way 10 way 5 way 10 way 5 way 10 way 4 way 6 way 4 way 6 way
+MsFNER 0.3442 0.3808 0.5616 1.0546 0.3482 0.3806 0.6708 1.2266 0.5126 0.4981 0.7530 0.6055
+ChatGPT 4.5581 5.4897 8.1842 9.3481 3.2633 8.0579 4.7755 5.3409 2.3078 2.8610 4.9470 7.2926
 (3) Firstly, we compute the representation $\hat{e}^{\prime}_{k}$ of each detected entity ${e}^{\prime}_{k}$ by Eq.6.
-Then, we input $\hat{e}^{\prime}_{k}$ into two independent modules: on one hand, we feed the $\hat{e}^{\prime}_{k}$ into KNN to obtain the predicted result $p_{knn}$ based on the $\mathcal{D}_{knn}$ Wang et al.
-(2022b), and on the other hand we pass the $\hat{e}^{\prime}_{k}$ into the finetuned best model for entity classification to produce its predicted type distribution $p_{soft}$ by Eq.5.
+Then, we input $\hat{e}^{\prime}_{k}$ into two independent modules: on one hand, we feed the $\hat{e}^{\prime}_{k}$ into KNN to obtain the predicted result $p_{knn}$ based on the $\mathcal{D}_{knn}$ [19], and on the other hand we pass the $\hat{e}^{\prime}_{k}$ into the finetuned best model for entity classification to produce its predicted type distribution $p_{soft}$ by Eq.9.
 (4) The final prediction type of the detected entity ${e}^{\prime}_{k}$ is jointly determined by KNN and the entity classification model:
 $p(y|e^{\prime}_{k})=\lambda p_{knn}(y|e^{\prime}_{k})+(1-\lambda)p_{soft}(y|e^{\prime}_{k})$ (12)
-where the $\lambda$ is a hyper-parameter that makes a balance between KNN and softmax.
-3 Experiments
-3.1 Experiments Setup
-The details can be found at A.1.
-3.2 Main results
-Table 1 illustrates the main results of different methods on FewNERD dataset.
-The results demonstrate that MsFNER significantly outperforms all the previous state-of-the-art approaches.
-MsFNER attains average F1 improvements of 2.65 and 4.44 on INTRA and INTER, respectively, compared to the previous best method, MAML-ProtoNet.
-In the 5-way 5-10 shot setting on the INTER dataset, we achieve the most substantial improvement over the MAML-ProtoNet method, reaching 7.79 points.
-Additionally, we perform the few-shot task employing ChatGPT, where MsFNER surpasses ChatGPT in F1 performance by an average score of 8.465 in the 5-10 shot setting.
-Furthermore, MsFNER surpasses ChatGPT by an average of 10.13 percentage points in the 1-2 shot setting in the INTER dataset.
-Overall, regardless of whether it is ChatGPT or small models, increasing the number of $K$-shot yields superior performance, while an increase in the number of $N$-way results in worse performance.
-3.3 Ablation Study
-To investigate the contributions of different modules of MsFNER, we conduct the ablation study by removing each of them at a time to observe the performance of our model.
-The results are shown at the bottom of Table 1.
-Firstly, we remove the contrastive learning module.
-In the 1 shot experiment of this setting, KNN is unnecessary as there is only one sample in the support dataset.
-We can see that the removal of contrastive learning results in a decrease of 0.905 average score, which indicates the significant impact and necessity of contrastive learning for enhancing representation.
-Secondly, we remove the KNN module and reserve the contrastive learning module.
-The removal of KNN drops 0.523 average score.
-We can see that the impact of contrastive learning and KNN increases with the number of $N$-way and $K$-shot.
-Removing either of them will lead to a significant decrease in the performance of MsFNER.
-3.4 Model Efficiency
-To evaluate the efficiency of MsFNER, we compute the average inference time of MsFNER and ChatGPT in different settings. More details can be found at A.3.
-4 Conclusion
-In this paper, we propose the MsFNER, aiming to improve the performance and efficiency of few-shot NER.
-The MsFNER can be achieved by three stages: training, finetuning and inferring.
-We evaluate MsFNER on the open FewNERD dataset, and results show the advance of MsFNER compared with previous few-shot NER methods and LLM.
-5 Limitations
-  • In practical, the effectiveness of Few-shot NER is seriously affected due to the data sparsity and data imbalance.
-  • The abundant external knowledge is very helpful for few-shot NER, which is worth exploring and studying.
-6 Ethical Considerations
-All the data we get is the open source, and there is no theft. We guarantee the originality, research ethics, societal impact and reproducibility of our work.
-Appendix A Appendix
-A.1 Experiments Setup
+where the $\lambda$ is a hyper-parameter that makes a balance between KNN and prototype classifier.
+Figure 2: The schematic diagram of the inference process for entity classification.
+III Experiments
+III-A Experiments Setup
 We conduct experiments on the Tesla V100 GPU.
-A.1.1 Datasets
-We conduct experiments on the FewNERD dataset Ding et al.
-(2021) which is annotated with 8 coarse-grained and 66 fine-grained entity types.
-We evaluate our method on the two settings of FewNERD: 1) FewNERD-INTRA, which pertains to a scenario wherein entities within the training set (source domain), validation set, and test set (target domain) are exclusively associated with distinct coarse-grained types.
-2) FewNERD-INTER, where only fine-grained entity types that do not intersect in different sets.
-Furthermore, we follow the settings of ESD Wang et al.
-(2022a) and adopt $N$-way $K\sim 2K$-shot sampling method to construct tasks.
-A.1.2 Parameter Settings
+III-A1 Datasets
+We conduct experiments on the FewNERD dataset [20] which is annotated with 8 coarse-grained and 66 fine-grained entity types.
+We evaluate our method on the two settings of FewNERD-INTRA and FewNERD-INTER, which are with the different grained types.
+Furthermore, we follow the settings of ESD [10] and adopt $N$-way $K\sim 2K$-shot sampling method to construct tasks.
+Another dataset we use for testing our method is FewAPTER, which is built by us based on the cybersecurity threat intelligence dataset APTER[21].
+Specifically, APTER is proposed by Wang et al.[21] within general data distributions, and it predefines 37 entity types under domain expert knowledge.
+In this paper, we refine and enhance the original APTER dataset, consolidating its entity types into 21 categories, resulting in a total of 28,250 entities.
+Inspired by the design methodology of the FewNERD dataset, we follow the knowledge transfer principle (from original entity type data $\rightarrow$ to new entity type data) and construct the FewAPTER dataset based on APTER for few-shot entity recognition research in the cybersecurity domain.
+Considering the data volume of each entity type, the 21 entity types are categorized into three subsets: the training set with 7 types {VULID, FILE, SECTEAM, LOC, TOOL, APT, ACT}, the validation set with 7 types {SHA1, OS, URL, MAL, DOM, PROT, TIME}, and the test set with 7 types {ENCR, VULNAME, IP, MD5, SHA2, IDTY, EMAIL}.
+While constructing the dataset, it is necessary that the entity types in the training, validation, and test sets do not overlap.
+Therefore, a maximum of 7-way can be set.
+However, due to the imbalanced data distribution and the possibility of multiple entity types occurring in a single sentence, setting the dataset to 7-way limits the diversity of knowledge learned, making it challenging to adapt to and generalize for new tasks.
+To address this, this paper constructs four $N$-way $K$-shot datasets: {4-way 1-shot, 4-way 3-shot, 6-way 1-shot, 6-way 3-shot}.
+Each $N$-way $K$-shot dataset includes 4000 training samples, along with 800 validation samples and 800 test samples.
+During the construction of each $N$-way $K$-shot dataset (for training, validation, and testing), each task, denoted as $S^{(i)}$ or $Q^{(i)}$, is randomly selected from the corresponding dataset, consisting of $N$ entity types and $K$ examples.
+For example, when constructing a task $S^{(i)}_{source}$ or $Q^{(i)}_{source}$ for the 6-way 3-shot training set, 6 entity types are randomly chosen from the 7 available types, with 3 entity examples selected for each entity type.
+III-A2 Parameter Settings
 We use grid search for hyperparameter settings and we train our model for 1,000 steps and choose the best model on validation dataset for testing.
@@
 We use the AdamW optimizer with a learning rate of 3e-5 and 0.01 linear warmup steps.
-The $K$ in KNN is set to 10 and the max-loss coefficient $\lambda$ of KNN is 0.1 when adding KNN.
+The $K$ in KNN is set to 10 and the coefficient $\lambda$ of KNN is 0.1 when adding KNN.
 The batch size is set to 32, the max sequence length is set to 128 and we use a dropout probability of 0.2.
-Models FewNERD-INTRA FewNERD-INTER
-1-2 shot 5-10 shot 1-2 shot 5-10 shot
-5 way 10 way 5 way 10 way 5 way 10 way 5 way 10 way
-MsFNER 0.3442 0.3808 0.5616 1.0546 0.3482 0.3806 0.6708 1.2266
-ChatGPT 4.5581 5.4897 8.1842 9.3481 3.2633 8.0579 4.7755 5.3409
-Table 2: The comparison of inference time between MsFNER and ChatGPT.
-A.1.3 Baselines
+III-B Baselines
 We compare MsFNER with popular baselines as follows:
-  • ProtoBERT Fritzler et al. (2019) employs the prediction of query labels by leveraging the similarity between the BERT hidden states derived from the support set and the query tokens.
-  • CONTAINER Das et al.
-(2022) employs token-level contrastive learning to train BERT as the token embedding function.
-Subsequently, BERT undergoes finetuning on the support set, followed by the application of a nearest neighbor method during the inference phase.
-  • NNShot Yang and Katiyar (2020) bears resemblance to ProtoBERT in its methodology, as it conducts predictions utilizing nearest neighbor algorithms.
-  • StructShot Yang and Katiyar (2020) incorporates an augmented Viterbi decoder into the inference process atop NNShot’s framework.
-  • ESD Wang et al. (2022a) formulates few-shot sequence labeling as a span-level matching problem and decomposes it into span-related procedures.
-  • MAML-ProtoNet Ma et al.
-(2022) proposes a decomposed meta-learning approach for few-shot NER, which tackles few-shot span detection and entity typing using MAML to facilitate rapid adaptation to novel entity classes and proposes MAML-ProtoNet for improved representation of entity spans.
-  • ChatGPT chatgpt (2023) is an advanced neural language model developed by OpenAI, which is trained on diverse corpora of text and could generate human-like text.
-A.2 ChatGPT Prompt Template
+  • ProtoBERT [5] employs the prediction of query labels by leveraging the similarity between the BERT hidden states derived from the support set and the query tokens.
+  • CONTAINER [9] employs BERT as the encoder and designs a token-level contrastive method to enhance the token embedding for few-shot NER.
+  • NNShot [8] bears resemblance to ProtoBERT in its methodology, as it conducts predictions utilizing nearest neighbor algorithms.
+  • StructShot [8] incorporates an augmented Viterbi decoder into the inference process atop NNShot’s framework.
+  • ESD [10] formulates few-shot sequence labeling as a span-level matching problem and decomposes it into span-related procedures.
+  • MAML-ProtoNet [12] proposes a meta-learning approach with fusing MAML and prototype network for few-shot NER, which facilitates rapid adaptation to novel entity types.
+  • BDCP [18] is a few-shot NER method with boundary discrimination and correlation purification, which explores the robustness of the model’s cross domain transfer learning ability in textual adversarial attack scenarios.
+  • ChatGPT [22] is an advanced large language model (LLM) developed by OpenAI, which is trained on diverse corpora of text and could generate human-like text.
+III-C ChatGPT Prompt Template
 In this section, we introduce the ChatGPT few-shot NER prompt template that we use. The prompt is composed of 3 parts: task description, few-shot cases, and input queries.
-In the task description, we initially elucidate the few-shot Named Entity Recognition (NER) task required from ChatGPT.
+In the task description, we initially elucidate the few-shot NER task required from ChatGPT.
 Subsequently, we provide ChatGPT with explicit instructions for implementing this task, including delineation of the entity types involved.
@@
 The template is as follows:
-Your task is to perform fewshot Named Entity Recognition. You could perform the task by the following actions:
+Your task is to perform few-shot Named Entity Recognition. You could perform the task by the following actions:
   1. 1.
@@
   3. 3.
-    Output a JSON object that contains the following keys: text, entity_list. Note that each item in entity_list is a dictionary with "entity" and "type" as keys.
+    Output a JSON object that contains the following keys: text, entity_list. Note that each item in entity_list is a dictionary with “entity” and “type” as keys.
 Here are the given few-shot cases: ‘‘‘\$case_input‘‘‘
@@
 The \$case_input comprises the sentences in the support set corresponding to $K$-shot, while the \$case_output represents the labels in the support set.
-Since ChatGPT is better at structured output, we use JSON format for output, with each output format being: {"entity": entity text ; "type": entity type}.
+Since ChatGPT is better at structured output, we use JSON format for output, with each output format being: {“entity”: entity text ; “type”: entity type}.
 The output will be merged into a list.
@@
 When we input this prompt to ChatGPT, it will return a list of relevant entity-type dictionaries.
-A.3 Model Efficiency
-To evaluate the efficiency of MsFNER, we compute the average inference time of MsFNER and ChatGPT in different settings.
-And the results are shown in Table 2, where we can see that MsFNER is on the millisecond level while the ChatGPT is on the second level.
-The time consumption is related to the amount of input data.
-In addition, the time consumption of ChatGPT will increase due to network impact.
-A.4 training process
-As the Figure 3 shows, the training loss on training dataset is smoothly declining within the same source types.
-But when finetuning on the validation set, there are some fluctuations due to the cross domain transfer of knowledge.
-And with the training time increasing, the overfitting is occurred slightly on validation dataset because of few supervised samples.
-Figure 3: The performance of training process on FewNERD 5-way 5~10-shot. (a) is the evaluation F1 score when finetuning on the validation dataset. (b) is the training loss on train dataset.
+III-D Main results
+Table I illustrates main results of different methods on FewNERD dataset and Table III shows results on the FewAPTER dataset.
+These results demonstrate that MsFNER significantly outperforms previous few-shot NER approaches.
+We have several findings: 1) MsFNER attains average F1 improvements of 2.65% and 4.44% on FewNERD-INTRA and FewNERD-INTER, respectively, compared to the previous best few-shot NER method, MAML-ProtoNet.
+Furthermore, the greater improvement is on the FewAPTER dataset, which achieves an average F1 score improvement of 11.42%.
+2) Although MsFNER outperforms the ChatGPT on the FewNERD dataset, it is slightly inferior to the LLM on FewAPTER.
+This is due to the characteristics of the FewAPTER dataset, which enable the model to exhibit strong learning capabilities within a specific domain, but with slightly reduced generalization ability for cross-domain learning.
+3) Regardless of whether it is ChatGPT or small models, increasing the number of $K$-shot yields superior performance, while an increase in the number of $N$-way results in worse performance.
+This aligns with general human intuition: given a fixed number of entity types, having more examples leads to better learning accuracy; conversely, with a fixed number of examples, increasing the number of entity types makes learning more challenging.
+4) From the performance evaluation of these results, it can be seen that the performance of LLM is quite stable, which significantly indicates that its influence from data and domain changes is relatively limited.
+This is primarily attributed to the large number of parameters in the model, which, unlike few-shot learning models, exhibits lower sensitivity to the number of training samples.
+III-E Ablation Study
+To investigate the contributions of different modules of MsFNER, we conduct the ablation study by removing each of them at a time.
+The results are shown at the bottom of Table I and Table III.
+Firstly, we remove the contrastive learning module.
+In the 1-shot experiment of this setting, KNN is unnecessary as there is only one sample in the support dataset.
+We can see that the removal of contrastive learning results in an average decrease of 0.905% on FewNERD and 0.745% on FewAPTER, which indicates the significant impact and necessity of contrastive learning for enhancing representation.
+Secondly, we remove the KNN module and reserve the contrastive learning module.
+The removal of KNN drops average 0.524% on FewNERD and 0.635% on FewAPTER.
+We can see that the impact of contrastive learning and KNN changes with the number of $N$-way and $K$-shot.
+Removing either of them will lead to a decrease in the performance of MsFNER.
+III-F Model Efficiency
+To comprehensively evaluate MsFNER and ChatGPT, we compute their average inference time(s) in different settings, and the results are shown in Table II.
+From the table, we can see that MsFNER is much faster than its LLM peer ChatGPT.
+According to the Occam’s Razor, we think that MsFNER is a better choice for few-shot NER in application.
+TABLE III:  F1 scores with standard deviations(%) on FewAPTER
+Models FewAPTER
+1 shot 3 shot
+4 way 6 way 4 way 6 way
+ESD† 33.83±0.57 29.97±0.42 41.60±0.27 39.08±0.72
+CONTAINER† 29.82 31.17 51.71 50.65
+BDCP† 44.65±0.97 36.55±1.02 51.43±0.63 39.55±1.74
+MAML-ProtoNet† 48.41±0.31 39.32±0.19 49.51±0.47 37.16±0.11
+ChatGPT‡ 59.14 57.65 61.14 57.68
+MsFNER 57.55±0.39 49.90±0.90 61.36±0.59 51.26±1.38
+w/o cl* 60.78±0.30 50.35±0.24
+w/o KNN 56.47±0.69 49.23±0.80 61.21±0.17 50.62±1.23
+III-G Training and Validation Process
+[…]
+Figure 3: The performance of training and validation stage on two different settings
+IV Conclusion
+In this paper, we propose MsFNER, a novel approach to addressing the challenges of a large number of negative samples and computational overhead in the domain of few-shot named entity recognition.
+MsFNER is implemented in three stages: training, fine-tuning, and inference.
+During the training and fine-tuning stages, we first detect the entity spans and subsequently recognize the entity types.
+To enhance the learning of entity representations, we introduce a contrastive learning module.
+In the inference stage, we replace the contrastive learning module with a KNN module to further improve inference performance.
+The final inference output is determined by combining the KNN results and the type classification predictions.
+We conduct experiments on the FewNERD and FewATPER datasets to evaluate the performance of MsFNER.
+Additionally, we compare our approach with ChatGPT-3.5 on few-shot tasks, and the results demonstrate that MsFNER can achieve performance comparable to ChatGPT-3.5 to some extent.